OpenStudy (cyclone3260):

Find the product. y^4 · y^6 Is it y^10??

OpenStudy (anonymous):

When you are multiplying exponents, they are adding them y^(4+6) = y^10 You are absolutely correct :)

OpenStudy (cyclone3260):

So then n^12 ÷ n^4 Would equal n^16?

OpenStudy (anonymous):

The rule is different for division. When you divide exponents, you actually subtract them = n^(12-4) = n^8 Does that make sense?

OpenStudy (cyclone3260):

Kay, so multiplying you add the exponents and dividing you subtract them makes sense! Any other rule like if I were to ADD, instead of multiplying? lol

OpenStudy (anonymous):

Actually there is one more rule. If you see something like (n^2)^4 you have to multiply the exponents. =n^(2)(4) = n^8

OpenStudy (anonymous):

That's: